OpenStudy (anonymous):

MEDAL REWARDED Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x4 - 5x2 - 36

OpenStudy (anonymous):

@John_ES

OpenStudy (anonymous):

you know it looks like \[(x+2i)(x-2i)(something)\] or \[(x^2+4)(something)\] find the "something" by division

OpenStudy (anonymous):

complex zeros always occur in pairs.

OpenStudy (anonymous):

or factor: \[ x^4 - 5x^2 - 36 =(x^2+4)(x^2-9)\]

OpenStudy (anonymous):

you didn't have to be told any zeros for this one to find all of them it is a quadratic in \(x^2\)

OpenStudy (anonymous):

i wanted to say if -2i is a zero then +2i is also a zero

OpenStudy (john_es):

I would follow @satellite73 indication, if -2i is a zero, then +2i is also a zero. Then, from this, I would divide, \[(x^4-5x^2+36)/(x^2+4)=x^2-9\] Then factorize it, \[x^2-9=(x+3)(x-3)\] And you find the other solutions, +3 and -3. In total, +2i, -2i, +3 and -3.

OpenStudy (john_es):

Sorry it should be -36 not +36.

OpenStudy (anonymous):

These are my choices: 2i, 6i, -6i 2i, 3i, -3i 2i, 6, -6 2i, 3, -3

OpenStudy (anonymous):

@John_ES

hero (hero):

@MandyNeedsHelp, @JohnES practically gave you the answer already.

OpenStudy (anonymous):

@Hero i dont want your help if you're just going to be impatient.

hero (hero):

I'm not being impatient. I'm just saying he made it more than obvious for you.